several books are on sale at a bookstore. some of these books cost 3 dollars , while other cost 5 dollars. one day last week, 80 of these books were sold. the total amount of sales was 300 dollars. how many of each type was sold?
1. write the equations that would solve the system
2show and explain how you solved the problem
3 how do you know you have the correct solution


Answer by stanbon(75887)

several books are on sale at a bookstore. some of these books cost 3 dollars , while other cost 5 dollars. one day last week, 80 of these books were sold. the total amount of sales was 300 dollars. how many of each type was sold?
1. write the equations that would solve the system
Quantity Equation: t + f = 80 books
Value Equation:: 3t +5f = 300 dollars
-----------------------------------------------
2show and explain how you solved the problem
Multiply thru the Quantity Eq. by 3 to get:
3t+3f = 3*80
Subtract that from the Value Eq. to get:
2f = 60
f = 30 (# of 5 dollar books sold)
Since t + f = 80, t = 50 (# of 3 dollar books sold)
------------------------------------------------------
3 how do you know you have the correct solution
Check the Value:
5*30+50*3 = 150+150 = $300
=============================
Cheers,
Stan H.